OpenStudy (anonymous):

please help :( What is the simplified form of the following expression? (x^10)(x^2) i will give best response and medal!

OpenStudy (anonymous):

x^20

OpenStudy (campbell_st):

the index law from multiplying the same base is add the powers \[x^a \times x^b = x^{a + b}\] you need to apply this you your question \[(x^{10})(x^{2}) = x^{10} \times x^2 \]